# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v4

@@ 3.3. Executing via Alternative Command Line
In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line. By inputting the algorithm list, problem list, and other parameters into the script file <mto.m>, experiments can be executed directly from the command line. For example, running the command mto({MFEA, EMEA}, {CMT1, CMT2}, 30) will execute the MFEA and EMEA algorithms on the CMT1 and CMT2 problems independently 30 times. Upon completion, the experiment data <MTOData.mat> is returned. Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
4. Extending MToP
Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP. By adhering to the code pattern outlined in Section 2.3, users can develop and implement their code files with ease.
4.1. Adding New Algorithms
To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the <Algorithms/> folder. This class file should inherit from the base class <Algorithm.m>. Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo. Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended. However, users can also choose to reimplement these functions to suit their specific requirements. Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
4.2. Adding New Problems
To integrate new problems into MToP, users should create a dedicated problem class file within the <Problems/> folder. This class file should inherit from the base class <Problem.m>. It’s advisable to define the suggested maximum function evaluations <maxFE> within the constructor of the specific problem. Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function. In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points. Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
4.3. Adding New Performance Metrics
To incorporate new performance metrics into MToP, users should create a new metrics function file within the <Metrics/> folder. Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object <MTOData> and return the calculated <result> object. For the metrics displayed in the GUI table, ensure that the <RowName>, <ColumnName>, and <TableData> fields are properly populated. Typically, the <RowName> for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task. The <ColumnName> typically represents the algorithm name. Implement the calculation function for the specific performance metric in the <TableData> field. For plotting metric convergence, ensure the <ConvergeData> field is populated. In the case of multi-objective optimization, include non-dominated solutions in the <ParetoData> field.
4.4. Extending Optimization Types
There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking (Tan et al., 2021). Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types. Users can refer to the existing code patterns in MToP to develop new optimization types. Specifically, for the sequential transfer optimization, it is possible to set the target task as the first task and the remaining tasks as source tasks for algorithms and problems. The metric calculation can be adjusted to evaluate the performance of the target task. For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
5. Conclusion and Future Work
@@
This paper introduces MToP, an open-source MATLAB platform designed for EMT. MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures. The current version of MToP contains over 50 MTEAs, more than 50 single-task EAs capable of handling MTO problems, over 200 benchmark MTO problems, and several real-world applications of EMT. This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
@@
While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some reproducing errors and bugs. Continuous efforts are underway to identify and rectify such issues, with updates regularly posted on GitHub. The platform encourages feedback and contributions from followers and researchers, with many codes already received and integrated from the community.
@@
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng et al., 2016; Deb and Jain, 2014), multitask combinatorial optimization codes (Feng et al., 2021c, a) and more real-world applications (Gupta et al., 2022; Huang et al., 2023) in the future.
@@ Acknowledgements.
This work was partly supported by the National Natural Science Foundation of China under Grant No. 62076225 and 62403477, the Open Research Fund of the State Key Lab of Digital Manufacturing Equipment & Technology under Grant No. IMETKF2024020, and the Major Project of Hubei Province Science and Technology under Grant No. 2023BCA006. All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP. The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian et al., 2017), for which the authors are also thankful.

### v5

@@ 4.4. Executing via Alternative Command Line
For users who require batch processing, integration with external scripts, or operation in non-GUI environments, MToP provides a comprehensive and flexible command-line interface. The main mto.m script has been created to serve as a dual-purpose entry point: calling mto with no arguments launches the GUI, while calling it with arguments executes a command-line experiment.
The function MTOData = mto(varargin) is designed for maximum flexibility, accepting both positional arguments and Name-Value pairs. For instance, users can run basic experiments using commands like result = mto(MFEA, CMT1) for quick runs. More complex configurations such as mto({MFEA, EMEA}, {CMT1, CMT2}, ’Reps’, 30, ’Par_Flag’, true, ’Save_Name’, ’A.mat’) allow for detailed customization, including parallel execution and specifying output filenames. Upon completion, the function returns the standard MTOData struct and simultaneously saves it to the specified .mat file. This command-line interface workflow extends beyond just running experiments; the returned MTOData object can be directly passed to any metric function to calculate results, just as in the GUI. This enables a complete, script-based workflow: setup, run, and analyze.
A comprehensive set of examples demonstrating this full process including how to instantiate and modify algorithm/problem objects, run experiments, calculate metrics like Obj and IGD, and programmatically plot the final convergence curves and Pareto fronts is provided in the cmd_examples.m script located in MToP’s root directory.
5. Validation and Comparison
This section first performs reproducible validation tests to ensure that the algorithms and problems implemented in MToP are reliable and accurate. Following this, a performance comparison between MToP and a popular EC platform, PlatEMO (Tian2017PlatEMO), is conducted to demonstrate MToP’s comparable efficiency.
5.1. Reproducible Validation
A core requirement for a scientific platform is the reliability and reproducibility of its components. To validate the implementations of the extensive algorithm and problem libraries within MToP, we conducted a comprehensive experimental study, running all algorithms (including both MTEAs and single-task EAs) on all compatible benchmark suites for 30 independent runs.
To ensure fair comparisons and high reproducibility, a global random seed controller is employed. For each independent repetition rep, the random number generator is explicitly seeded using rng(Global_Seed + rep - 1). This ensures that all algorithms within the same repetition begin with identical random conditions, which is crucial for reliable comparisons. In order to enhance reproducibility for the wider community and to provide a baseline for future research, we have made the complete raw experimental data from this large-scale validation publicly available. This dataset validates our implementations and can be used directly by other researchers to avoid redundant, computationally expensive experiments. All data can be accessed and downloaded. [^3] Pre-run experiment data can be found at:
  Google Drive: https://drive.google.com/drive/folders/1IpwXNuOYcnpC3IXbAx3VnLGVV899k3bG?usp=share_link
  Hugging Face: https://huggingface.co/datasets/intLyc/MToP-MTOData/tree/main
  Baidu Netdisk: https://pan.baidu.com/s/1Pk06fBj_4gidkiZe4f1Oww?pwd=mtop
5.2. Comparison with Other Platforms
Figure 12. Running time comparison between MToP and PlatEMO on DTLZ problems using algorithms: NSGA-II, SPEA2, MOEA-D, and SMS-EMOA.
Beyond internal validation, it is crucial to benchmark MToP’s performance and efficiency against established, state-of-the-art platforms. We selected PlatEMO (Tian2017PlatEMO), a widely-used and highly-regarded platform for evolutionary computation, as the basis for our comparison. To ensure a fair and direct comparison, we selected four popular and well-understood single-task multi-objective algorithms that are implemented in both platforms: NSGA-II, SPEA2, MOEA-D, and SMS-EMOA. We executed these algorithms on the standard DTLZ1-7 benchmark suite. Both platforms were run under identical experimental settings, including 30 independent runs and the use of parallel execution.
First, we compared the computational efficiency. Figure 12 displays the average wall-clock run time for each algorithm across the entire DTLZ suite. The results show that the execution times for MToP and PlatEMO are highly comparable. For NSGA-II, SPEA2, and MOEA-D, the run times are nearly identical. For SMS-EMOA, which involves computationally intensive hypervolume calculations, both platforms show similar overhead. This demonstrates that MToP’s architecture is lightweight and efficient, with no significant computational overhead compared to PlatEMO.
Second, we validated the correctness of our algorithmic implementations by comparing the final solution quality. Table LABEL:tab:results-dtlz presents the mean and standard deviation of the IGD metric for 30 runs. A Wilcoxon rank-sum test (at a 0.05 significance level) was conducted between the MToP and PlatEMO results for each case. The + / - / = summary shows that there is no statistically significant difference in the vast majority of cases (23 out of 28). The few minor differences are likely due to small implementation-level variations, but the overall performance is statistically indistinguishable. This similarity in performance is further confirmed visually in Figure LABEL:fig:pareto-comparison, which plots the final Pareto fronts from a sample run for each algorithm. The acquired fronts from both platforms are visually coincident.
6. Discussion and Outlook
@@
This paper introduces MToP, an open-source MATLAB platform designed for EMT. MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures. The current version of MToP contains over 50 MTEAs, more than 50 single-task EAs capable of handling MTO problems, over 200 benchmark MTO problems, and several real-world applications of EMT. This paper provides recommendations for selecting algorithms and benchmarks on different types of MTO problems. It also offers guidelines for using MToP, including testing algorithms and problems, executing comparative experiments, processing experiment data, and running via the command line. Finally, this paper validates the reproducibility of implemented algorithms and problems in MToP, provides open-sourced experimental data, and compares the running time and results of MToP with other EC platforms.
@@
While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some implementation errors and bugs. Continuous efforts are underway to identify and rectify such issues, with updates regularly posted on GitHub. MToP encourages feedback and contributions from users and researchers, with many codes already received and integrated from the community.
@@
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng2016RVEA; Deb2014NSGA3), multitask combinatorial optimization codes (Feng2021VRPHTO; Feng2021EEMTA) and more real-world applications (Gupta2021Half; Huang2023Evolutionary) in the future.
@@ Acknowledgements.
This work was partly supported by the National Natural Science Foundation of China (Grant No. 62076225), the Major Research & Development Program of Hubei Province (Grant No. 2025BEB002 & 2023BCA006), the Regional Innovation System Program of Hubei Province (Grant No. 2025EIA022), and the Hubei Superior and Distinctive Discipline Group of “New-Energy Vehicles and Smart Transportation”. All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP. The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian2017PlatEMO), for which the authors are also thankful.